# CRESSim–MPM: A Material Point Method Library for Surgical Soft Body Simulation with Cutting and Suturing

### v1

@@ I INTRODUCTION
  • A CUDA-based MPM library supporting multiple solvers with backend C-style APIs for efficient integration with external applications and other physics solvers.
@@
  • A preliminary evaluation of the simulation speed using different solvers and number of particles.
@@ II-A Surgical Simulators and Surgical Robotics
The second stream arises from the recent trend in surgical robotics research, particularly for surgical automation and autonomy, which relies on simulators for benchmarking, synthesizing data, and trial-and-error learning. Existing studies such as AMBF [13], ATAR [14], and V-Rep Simulator for the dVRK [15] provide diverse applications for surgical robotics research. Other work such as ORBIT-Surgical [16], Surgical Gym [17], FF-SRL [18] LapGym [19], AMBF-RL [20], SurRoL [21], dVRL [22], and UnityFlexML [1] has specific focus on machine learning applications. These platforms typically rely on an existing physics engine, such as PhysX (used in Nvidia Omniverse), FleX, SOFA, and Bullet. Therefore, their simulation capability largely depends on the physics solvers. For instance, Bullet is primarily a rigid body engine and only provides limited FEM soft body support. PhysX 5 supports FEM, but interactions such as cutting and suturing in soft bodies have not been realistically achieved, possibly due to the limitations of FEM. None of these studies uses MPM for soft body simulation, although it is particularly suitable for large deformation and fracture in surgical settings.
@@ III-A Standard MPM
$(m\mathbf{v})_{I}^{t+\Delta t}=\mathbf{f}_{I}^{t}(m\mathbf{v})_{I}^{t}\Delta t,\;\mathbf{v}_{I}^{t+\Delta t}=(m\mathbf{v})_{I}^{t+\Delta t}/m_{I},$ (3)
@@ IV-A Soft-Rigid Coupling
If the node is separating from the rigid body, the velocity is not changed, i.e., , $\mathbf{v}^{\prime}_{rel}=\mathbf{v}_{rel}$. Otherwise,
@@ VI SURGICAL SIMULATION EXAMPLES
This section includes screenshots from two example scenes representing surgical cutting and suturing. While the visuals and rendering are not highly realistic, the examples are primarily designed to showcase the physics capabilities. For demonstration purposes, we use two different solvers, although either one works in both scenarios.

### v2

@@ I INTRODUCTION
  • A CUDA-based MPM library supporting multiple MPM solvers with C-style APIs for efficient integration with external applications and other physics solvers.
@@
  • A preliminary evaluation of the simulation speed using different MPM solvers and number of particles.
@@ II-A Surgical Simulators and Surgical Robotics
The second stream arises from the recent trend in surgical robotics research, particularly for surgical automation and autonomy, which relies on simulators for benchmarking, synthesizing data, and trial-and-error learning. Existing studies such as AMBF [13], ATAR [14], and V-Rep Simulator for the dVRK [15] provide diverse applications for surgical robotics research. Other work such as ORBIT-Surgical [16], Surgical Gym [17], FF-SRL [18], LapGym [19], AMBF-RL [20], SurRoL [21], dVRL [22], and UnityFlexML [1] has specific focus on machine learning applications. These platforms typically rely on an existing physics engine, such as PhysX (used in Nvidia Omniverse), FleX, SOFA, and Bullet. Therefore, their simulation capability largely depends on the physics solvers. For instance, Bullet is primarily a rigid body engine and only provides limited FEM soft body support. PhysX 5 supports FEM, but interactions such as cutting and suturing in soft bodies have not been realistically achieved, possibly due to the limitations of FEM. A Recent study [23] integrates basic MPM support into SurRoL using MLS-MPM but does not address cutting and suturing contact methods. In contrast, our work focuses on developing a comprehensive engine-level MPM library specifically designed for simulating surgical soft body manipulations.
@@ III-A Standard MPM
$(m\mathbf{v})_{I}^{t+\Delta t}=(m\mathbf{v})_{I}^{t}+\mathbf{f}_{I}^{t}\Delta t,\;\mathbf{v}_{I}^{t+\Delta t}=(m\mathbf{v})_{I}^{t+\Delta t}/m_{I}^{t},$ (3)
@@ IV-A Soft-Rigid Coupling
If the node is separating from the rigid body, the velocity is not changed, i.e., $\mathbf{v}^{\prime}_{rel}=\mathbf{v}_{rel}$. Otherwise,
@@ VI SURGICAL SIMULATION EXAMPLES
This section includes screenshots from two example scenes representing surgical cutting and suturing. While the visuals and rendering are not highly realistic, the examples are primarily designed to showcase the physics capabilities. For demonstration purposes, we use two different solvers, although either one works in both scenarios. Additional examples are provided in the supplementary video.